When two people were talking in a small room, the sound intensity level was 61 dB everywhere within the room. Then, there were 10 people talking in similar manner simultaneously in the room, what was the resulting sound intensity level (in dB)?
Answer #1

Let, intensity of the sound of one people be I, in W / m2.

So, intensity of the sound of two people is 2I, in W / m2.

Hence, intensity of the sound of two people in dB scale is : 10 log ( 2I / Io ) = 61 dB, Io = 10-12 W / m2.

So, intensity of the sound of 10 people in dB scale is :

10 log ( 10I / Io ) = 10 log ( 5 x 2I / Io ) = 10 log 5 + 10 log ( 2I / Io ) = ( 6.9897 + 61 ) db ~ 68 dB.

Hence, when there were 10 people talking in similar manner simultaneously in the room, the resulting sound intensity level in the room was 68 dB.